A consultant traveled 4 hours to attend a meeting. The return trip took only 3 hours because the speed was 8 miles per hour faster. What was the consultant's speed each way?
Answer by checkley71(8403)

d=rt or in this case the distances are the same.
(x+8)3=4x
3x+24=4x
3x-4x=-24
-x=-24
x=24 mph on the first trip.
24+8=32 mph for the return trip.
proof:
32*3=4*24
96=96